CAN YOU PLEASE SHOW ME HOW TO SOLVE THIS PROBLEM .(8+5)+2=2+(n+8)
Answer by roxy9229(32)

(8+5)+2=2+(n+8)
13+2 = 2+n+8
15 = 10+n
subtract 10 from both sides
5=n
n=5